# Zero-Shot Position Debiasing for Large Language Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
LLM have demonstrated remarkable unsupervised ability in various tasks (Kojima et al., 2022). Even so, fine-tuning still overtakes it under the task-specific setting (Ding et al., 2023). However, fine-tuned LLMs might rely on the dataset biases and artifacts as shortcuts for prediction, as the fine-tuning datasets are sometimes skewed due to budget constraints (Du et al., 2022). Consequently, these LLMs have poor generalization performance when applied to unseen test data and are vulnerable to various types of adversarial attacks (Meade et al., 2022).
@@
Position bias has been shown in LLMs after fine-tuning (Liu et al., 2023). Specially, the well-known LLMs, e.g., GPT-3.5111https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates, longchat-13B [^2] https://lmsys.org/blog/2023-06-29-longchat, are skilled when the relevant information occurs at the beginning or end of the input context, while the performance significantly degrades when LLMs need to find relevant information in the middle of the context. Analysis of conversational question answering (CQA) on CANARD (Elgohary et al., 2019) dataset further confirms the existence of position bias. As shown in Fig. 1, after fine-tuning, T5-FT achieves about 0.8 in terms of ROUGE-L when relative position is 1, but less than 0.1 when relative position is 4. The striking contrast in performance attracts increasing attention from researchers on debiasing or mitigating position bias Meade et al. (2022).
@@
Figure 1: Question answering performance of T5-large (T5) and fine-tuned T5-large (FT) over different relative positions in CANARD. Relative position means the distance between utterances containing the last turn answer span and the current turn answer span.
@@
Early works mainly focus on mitigating position bias on extractive question answering (QA) before the emergence of LLMs (Ko et al., 2020; Karimi Mahabadi et al., 2020). A prominent debiasing method is Product-of-Expert (PoE), which discourages the extractive model from learning position bias picked up by the fixed biased model (Du et al., 2021; Shinoda et al., 2022). Recently, more and more works focus on generative language model debiasing for LLMs (Guo et al., 2022; Li et al., 2023). These works mostly focus on mitigating social bias (Kasneci et al., 2023), e.g., gender bias, and racial bias. They adopt either in-context learning (ICL) to guide the generation of LLM (Meade et al., 2023) or prompt tuning (PT) to fine-tune prompts for LLMs (Li et al., 2023). However, these works cannot be transferred to mitigate position bias, as they require manually annotated non-biased samples for ICL or external bias knowledge for PT, which are lacking for position bias.
In this paper, we propose to mitigate position bias for LLMs in zero-shot setting, i.e., without external bias knowledge or any non-biased samples. To achieve this, we propose to leverage the low position bias characteristics of pre-trained LLMs. Previous studies have shown that pre-trained LLMs are more robust to position bias (Utama et al., 2021). After being pre-trained on a large volume of data, pre-trained LLMs can leverage utterances from various positions in the document for generation, thus having low position bias. As shown in Fig. 1, pre-trained T5 achieves ROUGE-L score (Lin, 2004) ranging from 0.2 to 0.4 across almost all relative positions, demonstrating its robustness against position bias.
@@
Motivated by this, we propose a ZOE framework to mitigate position bias for LLMs. First, we use a low-bias inference module to collect unsupervised responses with low position bias by applying diverse prompting strategies. Then, we propose a MSA module to prune the unsupervised responses, as low-quality responses will also undermine model performance on non-biased samples. Finally, we use a multi-objective optimization module to additionally optimize with unsupervised responses. The whole process does not require any external bias knowledge or non-biased samples, which is zero-shot and general enough.
@@
To verify the effectiveness of ZOE, we have done experiments on five tasks with eight datasets. Experimental results show that the ZOE achieves superior performance in mitigating three types of position bias significantly, including lead bias, relative position bias and lexical bias. To summarize,
@@
  • We propose to mitigate position bias in a zero-shot setting, i.e., without any external knowledge or datasets.
@@
  • We propose a ZOE framework with a MSA module to prune low-quality unsupervised responses for fine-tuning.
@@
  • Experiments show that ZOE mitigates various types of position biases by sacrificing only small performance on biased samples, demonstrating its effectiveness and generalization.
@@ 2.1 Task Definition
Given a biased dataset $D$ for the target task, the task aims to fine-tune the pre-trained model to improve the robustness against position bias, i.e., to achieve superior performance on non-biased samples by retaining the performance on biased samples. Here, the biased samples exhibit position bias consistent with that of the training samples and the non-biased samples lie outside the distribution of the training samples. As the positions of useful utterances vary in real-world scenarios, it is more applicable to achieve improved performance on non-biased samples by sacrificing partial performance on biased samples.
[…]
Figure 2: Overview of our proposed zero-shot position debiasing (ZOE) framework (taking CQG as the example). It consists of three components: low-bias inference to collect unsupervised responses, master-slave alignment module to align unsupervised responses, multi-objective optimization to fine-tune with unsupervised responses.
@@ 2.2 Large Language Model
LLM have attracted much attention and become state-of-the-art due to their remarkable ability of language generation. They formulate all natural language processing (NLP) tasks as language generation tasks with different task prompts:
@@
where $x$, $y$ and $prompt$ are task input, output, and task prompt, respectively. And $y_{i}$ means the $i$-th token in $y$. Task prompt $prompt$ consists of the task instruction and demonstrations to tell the LLMs the definition of the task and how it works. The introduction of task prompt enables LLMs to utilize all available data for training and improve their generalization ability on unseen tasks.
3 Model
We propose a zero-shot position debiasing (ZOE) framework to mitigate position bias for generative LLMs. As shown in Fig. 2, ZOE consists of three parts: low-bias inference, master-slave alignment (MSA) and multi-objective optimization, where all modules do not require external bias knowledge or non-biased dataset. Low-bias inference generates unsupervised responses with lower position bias based on pre-trained LLMs (in §3.1). Then we prune these low-quality unsupervised responses based on the target responses by master-slave alignment (in §3.2). Finally, we fine-tune the LLMs by optimizing two objectives: target responses and aligned unsupervised responses (in §3.3).
@@ 3.1 Low-bias Inference
Given a biased training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ for the target task, the low-bias inference module aims to generate unsupervised low-bias responses based on pre-trained LLMs, which are more robust and have low position bias Utama et al. (2021).
We generate unsupervised responses through three prompting strategies and apply different prompting strategies based on the task type.
  • Instruction-only prompting generate responses of target task by feeding the task input and task instruction to the pre-trained LLMs. Concretely, we assign the $prompt$ by task instruction in Eq. 1 for the generation. We adopt this prompting strategy for three fact generation tasks, CQA, KGC and summarization, as their target responses are always unique and unambiguous in semantics.
@@
[…]
  • Diverse prompting obtains multiple responses of the target task by feeding various prompts to LLMs. It can be used in CQG, as various prompts can specify diverse aspects of the generated responses. For instance, we can specify a ‘What’ question for generation by feeding “Please generate a ‘What’ question based on the following dialogue context and passage”.
3.2 Master-slave Alignment
Given the generated unsupervised response $y^{\prime}_{i}$ for input $x_{i}$, and the target response $y_{i}$, the master-slave alignment (MSA) module prunes the unsupervised responses to align with the target response. The reason is that unsupervised responses are low-quality, as they are generated from pre-trained LLMs without training on the target task, which only serves as a slave objective for mitigating position bias. In contrast, the target responses are human-annotated with high quality, which is the primary or master objective for optimization.
Alignment for NLI. As the generated responses for NLI are deterministic, i.e., entailment, neutral and contradiction, we first estimate the probability distribution over all classes by prompting and then align it by masking. As maximizing the probability of the target class will strengthen the position bias of the training dataset, we align the probability distribution by masking the target class to mitigate position bias:
@@ 3.2 Master-slave Alignment
where $p(class_{j}|x_{i})$ is the generation probabilities of tokens for $j$-th class, which is calculated by Eq. 1.
Alignment for other tasks. We first identify low-quality unsupervised responses and then drop them, as modifying after identifying responses will lead to error accumulation. Four types of low-quality responses are identified in CQA, summarization, KGC and CQG: non-compliant responses, dull responses, incoherent responses and unreliable responses. We apply and combine four strategies to identify them.
  • Non-compliant identification identifies whether the unsupervised responses follow the task instruction by keyword matching. For example, it checks whether the generated question is a ‘What’ question when specified in the instruction.
  • Dull identification identifies whether the unsupervised responses are dull by keyword matching, e.g., “What is the title of the passage?”.
  • Incoherent identification identifies whether the unsupervised responses are incoherent based on perplexity. The response is dropped if the perplexity of any token in it falls below a pre-defined threshold.
  • Unreliable identification identifies whether the unsupervised responses are reliable by comparing them with the target response. The intuition is that the fact in the response may be wrong if its semantics deviate significantly from the fact in the reference. We identify and drop the response if the overlap score between it and the target response is less than a pre-defined threshold.
For CQA, summarization and KGC tasks, we align unreliable responses as the facts in answers, summaries and knowledge-enriched responses are always unique in semantics. For CQG task, we align non-compliant, dull or incoherent responses as the appropriate questions are diverse in semantics.
@@ 3.3 Multi-Objective Optimization
Given the target response $y_{i}$, aligned unsupervised response $y^{\prime}_{i}$ and input $x_{i}$, the multi-objective optimization module fine-tunes the model to generate high-quality but low-bias responses. It fine-tunes the model by optimizing two objectives: target responses as a master objective to improve the generation quality and aligned unsupervised responses as a slave objective to mitigate position bias:
@@
where $\alpha$ is a hyper-parameter and $\mathcal{L}_{target}(x_{i},y_{i})$ is a negative log-likelihood (NLL) loss to maximize the generation probability of target response:
@@
For NLI task, we maximize the probability of the most likely class after masking in Eq. 2:
@@
where ${ind(i)}$ is the index of the most likely class:
@@
$y^{\prime}_{i}$ is the masked probability distribution for all classes in Eq. 2, $s_{i,ind(i)}$ is its generation probability in Eq. 4 and $class_{ind(i)}$ is the corresponding tokens.
For other tasks, we use NLL loss to maximize of generation probability of $y^{\prime}_{i}$ for CQA and CQG, summarization and KGC:
@@ 4 Experiments
We evaluate ZOE on three categories of NLP tasks based on changes in conveyed information from input to output: language understanding tasks, language compression tasks and language creation tasks Deng et al. (2021). Language understanding tasks (e.g., NLI, CQA) aim to comprehend and interpret natural language input given a conversation or document context. For a compression task (e.g., summarization), the goal is to concisely describe the most important information in the input (e.g., a document). A creation task (e.g., CQG, KGC) generates output that adds new information on top of input (e.g., dialogue history).

### v3

# Self-Supervised Position Debiasing for Large Language Models

@@ Abstract
Fine-tuning has been demonstrated to be an effective method to improve the domain performance of large language models. However, LLMs might fit the dataset bias and shortcuts for prediction, leading to poor generation performance. Previous works have proven that LLMs are prone to exhibit position bias, i.e., leveraging information positioned at the beginning or end, or specific positional cues within the input. Existing debiasing methods for LLMs require external bias knowledge or annotated non-biased samples, which is lacking for position debiasing and impractical in reality. In this work, we propose a self-supervised position debiasing (SOD) framework to mitigate position bias for LLMs. SOD leverages unsupervised responses from pre-trained LLMs for debiasing without relying on any external knowledge. To improve the quality of unsupervised responses, we propose an objective alignment (OAM) module to prune these responses. Experiments on eight datasets and five tasks show that SOD consistently outperforms existing methods in mitigating three types of position biases. Besides, SOD achieves this by sacrificing only a small performance on biased samples, which is general and effective. To facilitate the reproducibility of the results, we share the code of all methods and datasets on https://github.com/LZKSKY/SOD.
Self-Supervised Position Debiasing for Large Language Models
Zhongkun Liu1,  Zheng Chen1,  Mengqi Zhang1,  Zhaochun Ren2,  Zhumin Chen1,   Pengjie Ren1thanks:  Corresponding authors. 1School of Computer Science and Technology, Shandong University, China 2 Leiden University, Leiden, the Netherlands {liuzhongkun,chenzheng01}@mail.sdu.edu.cn, {mengqi.zhang,chenzhumin,renpengjie}@sdu.edu.cn,  z.ren@liacs.leidenuniv.nl
@@ 1 Introduction
Although LLMs have demonstrated remarkable unsupervised ability in various tasks (Kojima et al., 2022), fine-tuning still overtakes it under the task-specific setting (Ding et al., 2023). However, fine-tuned LLMs might rely on the dataset biases and artifacts as shortcuts for prediction, as the fine-tuning datasets are sometimes skewed due to budget constraints (Du et al., 2022). This results in poor generalization performance when applying fine-tuned LLMs to unseen test data and these models are vulnerable to various types of adversarial attacks (Meade et al., 2022).
@@
Position bias has been demonstrated to exist across various fine-tuned LLMs (Liu et al., 2023). Specially, the well-known LLMs, e.g., GPT-3.5111https://openai.com/blog/gpt-3-5-turbo-fine-tuning-and-api-updates, longchat-13B [^2] https://lmsys.org/blog/2023-06-29-longchat, are skilled when the relevant information occurs at the beginning or end of the input context, while the performance significantly degrades when LLMs need to find relevant information in the middle of the context. Analysis of conversational question answering (CQA) on CANARD (Elgohary et al., 2019) dataset further confirms the existence of position bias. As shown in Fig. 1, 80% of the performance improvement after fine-tuning is attributed to fitting bias on relative positions 0-2. This encourages researchers to engage in position debiasing Meade et al. (2022).
@@
Figure 1: Question answering performance of FlanT5-large (T5) and fine-tuned FlanT5-large (FT) over different relative positions in CANARD. Relative position means the distance of grounded utterances between the last turn answer and the current turn answer.
@@
Early works mainly focus on mitigating position bias on extractive tasks before the emergence of LLMs (Ko et al., 2020; Karimi Mahabadi et al., 2020). A prominent debiasing method is Product-of-Expert (PoE), which discourages the extractive model from learning position bias picked up by the fixed biased model (Du et al., 2021; Shinoda et al., 2022). Recently, more and more works focus on debiasing for generative models, e.g., LLMs (Guo et al., 2022; Li et al., 2023). They adopt either in-context learning (ICL) to guide the generation of LLM (Meade et al., 2023) or prompt tuning (PT) to fine-tune prompts for LLMs (Li et al., 2023). However, these works mostly focus on mitigating social bias (Kasneci et al., 2023), e.g., gender bias and racial bias, leaving position debiasing unexplored. Besides, these works cannot be transferred to mitigate position bias, as they require manually annotated non-biased samples for ICL or external bias knowledge for PT, which are lacking for position debiasing.
To deal with this challenge, we propose to leverage the low position bias characteristics of pre-trained LLMs. Previous studies have shown that pre-trained LLMs are more robust to position bias (Utama et al., 2021). This is due to the randomness of knowledge utilization in generation during pre-training.
As shown in Fig. 1, the ROUGE-L score of pre-trained T5 fluctuates within the range of 0.2 to 0.4 across almost all relative positions, demonstrating its robustness against position bias.
@@
In this paper, we propose a SOD framework to mitigate position bias for LLMs. First, we use a low-bias inference module to collect unsupervised responses with low position bias by applying various prompting strategies. Then, we propose an OAM module to prune the unsupervised responses, as low-quality responses will undermine model performance on non-biased samples. Finally, we use a multi-objective optimization module to leverage these unsupervised responses for fine-tuning. The whole process does not require any external bias knowledge or non-biased samples, which is self-supervised and general.
@@
To verify the effectiveness of SOD, we conduct experiments on eight datasets covering five tasks. Experimental results show that the SOD achieves superior performance in mitigating three types of position bias significantly, including lead bias, relative position bias, and lexical bias. The main contributions of this work are as follows.
@@
  • We propose to mitigate position bias for LLMs in a self-supervised setting, i.e., without any external knowledge or annotated samples.
@@
  • We propose a SOD framework for position debiasing with an OAM module to prune low-quality unsupervised responses for fine-tuning.
@@
  • Experiments show that SOD can mitigate various types of position biases by sacrificing only small performance on biased samples, demonstrating its effectiveness and generality.
@@ 2.1 Task Definition
Given a biased dataset $D$ for the target task, our position debiasing task aims to improve the model robustness against position bias when fine-tuning on target tasks, i.e., to achieve superior performance on non-biased samples by retaining the performance on biased samples. Here, the biased samples exhibit similar position bias as training samples, and the non-biased samples do not contain these position clues. The target task for fine-tuning can be any natural language processing (NLP) tasks exhibiting position bias. In this paper, we focus on five target tasks: conversational question answering (CQA), conversational question generation (CQG), knowledge-based conversation (KGC), summarization and natural language inference (NLI).
Figure 2: Overview of our proposed self-supervised position debiasing (SOD) framework (taking CQG as the example). First, the low-bias inference module collects multiple unsupervised questions from LLMs. Then, the objective alignment module aligns these questions with the target question. Finally, these aligned questions are utilized for fine-tuning within the multi-objective optimization module.
@@ 2.2 Large Language Model
LLM have attracted much attention and become state-of-the-art due to their remarkable ability of language generation. They formulate all NLP tasks as language generation tasks with different task prompts:
@@
where $x$, $y$ and $prompt$ are task input, output, and task prompt, respectively. $y_{i}$ denotes the $i$-th token in $y$. Task prompt $prompt$ consists of the task instruction and demonstrations to tell the LLMs the definition of the task and how it works. The introduction of task prompt enables LLMs to utilize all available data for training and improve their generalization ability on unseen tasks.
3 Method
We propose a self-supervised position debiasing (SOD) framework to mitigate position bias for generative LLMs. As shown in Fig. 2, SOD consists of three modules: low-bias inference, objective alignment (OAM), and multi-objective optimization, where all modules do not require external bias knowledge or non-biased samples. The Low-bias inference module generates unsupervised responses with lower position bias by utilizing pre-trained LLMs (in §3.1). Subsequently, the OAM module is employed to prune these low-quality unsupervised responses based on the target responses (in §3.2). Finally, the multi-objective optimization module fine-tunes the LLMs by optimizing the task objective and the debiasing objective (in §3.3). The task objective utilizes target responses to enhance task-specific performance, and the debiasing objective leverages unsupervised responses for position debiasing.
@@ 3.1 Low-bias Inference
Given a biased training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ for the target task, the low-bias inference module generates unsupervised responses with low position bias based on pre-trained LLMs Utama et al. (2021).
We employ three prompting strategies for generation and adapt them for different target tasks.
  • Instruction-only prompting generates respon-
    ses of target task by feeding the task input and task instruction to the pre-trained LLMs. Concretely, we assign the $prompt$ by task instruction in Eq. 1 for the generation.
  • Diverse prompting generates responses with diverse aspects by feeding various prompts to LLMs.
@@
    ICL also feeds multiple input-output examples to LLMs for generation, in addition to the task instruction and input. It enhances the model comprehension of target tasks but requires a longer input length.
We adopt ICL only for NLI, due to the limit of the model input length. We employ diverse prompting for CQG, which is intrinsically creative and diverse. And instruction-only prompting is implemented for CQA, KGC, and summarization.
3.2 Objective Alignment
Unlike the annotated high-quality target responses for the task objective, the unsupervised responses for the debiasing objective are of lower quality and noisy, because they are generated by pre-trained models without specific fine-tuning on the target task. To reduce the interference between the debiasing objective and the task objective, we propose an objective alignment (OAM) module to prune the unsupervised responses $y^{\prime}_{i}$ to better align with the target response $y_{i}$.
We propose various alignment strategies for different tasks based on their intrinsic characteristics.
Alignment for tasks excluding NLI. We first identify low-quality unsupervised responses and then drop them, as the generated responses are flexible for modification and modification may introduce new errors. We apply and combine four strategies to identify them.
  • Non-compliant identification identifies unsupervised responses deviating from the task instruction by keyword matching. For example, it identifies non-‘what’ questions when ‘what’ is specified in the instruction.
  • Dull identification identifies dull responses by keyword matching, e.g., “What is the title of the passage?”.
  • Incoherent identification identifies incoherent responses if the perplexity of any token in the response falls below a pre-defined threshold.
  • Unreliable identification identifies unreliable responses if the overlap score between unsupervised and target responses is less than a pre-defined threshold. The intuition is that the fact in the response may be wrong if its semantics deviate significantly from the fact in the reference.
For CQA, summarization, and KGC tasks, we align unreliable responses since the facts in answers, summaries, and knowledge-enriched responses are always unique in their semantics. For CQG task, we align non-compliant, dull, or incoherent responses, considering that the appropriate questions are diverse in semantics.
Alignment for NLI. As the generated responses for NLI are deterministic, i.e., entailment, neutral, and contradiction, we can directly estimate the probability distribution over all classes by prompting and then align it. The estimated probability distributions are low-quality sometimes when the target class dominates, which is redundant for optimizing the task objective and strengthens position bias. Therefore, we align the estimated probability distribution by masking the target class:
@@ 3.2 Objective Alignment
where $p(class_{j}|x_{i})$ is the generation probabilities of $j$-th class tokens, which is calculated by Eq. 1.
@@ 3.3 Multi-Objective Optimization
Given the target response $y_{i}$, aligned unsupervised response $y^{\prime}_{i}$, and input $x_{i}$, our multi-objective optimization module fine-tunes the model to generate task-specific but low-bias responses. It fine-tunes the model by optimizing two objectives: target responses as task objective to improve the performance on the target task and unsupervised responses as debiasing objective to mitigate position bias:
@@
where $\alpha$ is a hyper-parameter for tradeoff and $\mathcal{L}_{target}(x_{i},y_{i})$ is a negative log-likelihood (NLL) loss to maximize the generation probability of target response:
@@
For tasks excluding NLI, such as CQA and CQG, summarization and KGC, we use NLL loss to maximize the probability of generating $y^{\prime}_{i}$:
For NLI, we maximize the generation probability of the most likely class tokens after alignment:
@@
where ${ind(i)}$ is the index of the class, $class_{ind(i)}$ is the class tokens, and $s_{i,ind(i)}$ is the generation probability of the class tokens in Eq. 4:
@@
$y^{\prime}_{i}$ is the masked probability distribution for all classes in Eq. 2.
@@ 4 Experiments
We evaluate SOD on three categories of NLP tasks based on changes in conveyed information from input to output: language understanding tasks, language compression tasks and language creation tasks Deng et al. (2021). Language understanding tasks (e.g., NLI, CQA) aim to comprehend and interpret natural language input given a conversation or document context. For a compression task (e.g., summarization), the goal is to concisely describe the most important information in the input (e.g., a document). A creation task (e.g., CQG, KGC) generates output that adds new information on top of input (e.g., dialogue history).